### v1

# A Frank-Wolfe-based primal heuristic for quadratic mixed-integer optimization

Abstract
We propose a primal heuristic for quadratic mixed-integer problems. Our method extends the Boscia framework – originally a mixed-integer convex solver leveraging a Frank-Wolfe-based branch-and-bound approach – to address nonconvex quadratic objective and constraints. We reformulate nonlinear constraints, introduce preprocessing steps, and a suite of heuristics including rounding strategies, gradient-guided selection, and large neighborhood search techniques that exploit integer-feasible vertices generated during the Frank-Wolfe iterations. Computational results demonstrate the effectiveness of our method in solving challenging MIQCQPs, achieving improvements on QPLIB instances within minutes and winning first place in the Land-Doig MIP Computational Competition 2025.
1 Introduction
Mixed-Integer Quadratically Constrained Quadratic Problems (MIQCQPs) represent a broad category in optimization, capturing an array of applications in operations research and machine learning. These problems combine the combinatorial difficulties of mixed-integer structures with nonlinear nonconvex constraints. Despite this double difficulty, solution methods, algorithms, and solvers have been developed over the past decades for generic or specific forms of MIQCQPs, including Couenne [2], GloMIQO [28], ANTIGONE [29], BARON [34], and SCIP [8, 10, 9]. We refer the reader to [24, 25] for recent reviews of mixed-integer nonlinear optimization beyond quadratic functions. Recently, commercial mixed-integer linear solvers such as Xpress and Gurobi added capabilities to solve MIQCQPs to global optimality, showing the strong interest from application areas. We denote MIQCQPs in the following form:
$\min_{x\in\mathbb{R}^{n}}\quad$ $\frac{1}{2}x^{\top}Qx+d^{\top}x$ (1)
[…]
$x^{L}_{k}\leqslant\ x_{k}\leqslant\ x^{U}_{k}$ $\quad\forall\,k\in N,$
$x_{k}\in\ \mathbb{Z}$ $\quad\forall\,k\in I,$
where $N:=\{1,\dots,n\}$, $I\subseteq N$ is the index set of integer variables, $Q,A_{i}\in\mathbb{Q}^{n\times n}$ are symmetric matrices, $d,\ b_{i}\in\mathbb{Q}^{n}$, $c_{i}\in\mathbb{Q}$ for $i=1,\dots,m$, and $x^{L},x^{U}\in\bar{\mathbb{Q}}^{n}$ (with $\bar{\mathbb{Q}}:=\mathbb{Q}\cup\{\pm\infty\}$) are the lower and upper bounds of $x$. Note that both $Q$ and $A_{i}$ need not be positive semidefinite, hence nonconvex quadratic problems are allowed. If $I=\emptyset$, the problem reduces to a quadratically constrained quadratic program (QCQP). In the following, when discussing a single quadratic constraint, we may omit the constraint index $i$ for simplicity.
In this work, we present a framework for this class of problems with an emphasis on finding primal feasible solutions, which we developed in the context of the Mixed-Integer Programming workshop 2025  [^1] Details on the competition can be found at the following page:
https://www.mixedinteger.org/2025/competition/
. The framework also provides dual bounds for a subset of MIQCQPs as a by-product of the primal search. Our approach is based on a branch-and-bound algorithm leveraging Frank-Wolfe to optimize continuous relaxations over the convex hull of mixed-integer feasible points, as proposed in the Boscia framework [21]. Importantly, our method can leverage mixed-integer linear optimization solvers to compute vertices of a relaxed feasible region and can naturally handle differentiable nonlinear terms in the objective function, a property we leverage to transform and relax constraints. Guided by the nonlinear relaxations, we fix continuous and integer variables and use other off-the-shelf solvers for the resulting small problems, adapting techniques from the large neighborhood search literature. A high-level overview of our framework is presented in Figure 1.
Figure 1: Overview of our approach. The model first undergoes a pre-processing whose main goal is to transfer nonlinearities into the objective function. Our workhorse is indeed the Frank-Wolfe-based solver Boscia which handles nonlinear problems (NLP) by suitably combining calls to a mixed-integer programming (MIP) solver in a branch-and-bound (BnB) framework. The points collected at each node in this process are then fed to various heuristics to reach better solutions.
The rest of the paper is organized as follows. Section 2 presents the core part of the framework consisting of the specialized branch-and-bound with relaxations solved with Frank-Wolfe. Section 3 details the transformations we operate on the original MIQCQP, modifying the set of constraints and the objective. Section 4 expands on the heuristics called during and after the branch-and-bound process. Section 5 presents computational results assessing the performance of the framework and the impact of its different components.
2 Boscia and Frank-Wolfe
Boscia [22] is a mixed-integer convex solver that tackles problems of the form
$\min_{x\in\mathbb{R}^{n}}\quad$ $f(x)$ (2)
s.t. $x\in\mathcal{X}$
$x_{i}\in\mathbb{Z}\quad\forall\,i\in I$
where $\mathcal{X}\subseteq\mathbb{R}^{n}$ is a convex, compact (polyhedral) set and $I\subseteq N$ is the set of integer variables.
It operates a branch-and-bound scheme, utilizing Frank-Wolfe (FW) methods [11] implemented in the FrankWolfe.jl library [6, 7] to solve continuous subproblems. FW methods are projection-free first-order methods that only require oracle access to the objective function, its gradient and the feasible region. At iteration $t$, the classic FW algorithm, dubbed Vanilla Frank-Wolfe, calls the Linear Minimization Oracle (LMO) over the feasible region using the gradient of the current iterate $x_{t}$ as the objective function. The solution $v_{t}$ is used to generate the next descent direction $d_{t}=x_{t}-v_{t}$ and the iterate is updated as $x_{t+1}=x_{t}-\gamma_{t}d_{t}$ where $\gamma_{t}$ is the step-size. Thus, the iterate can be represented as a convex composition of the extreme points of the feasible region. The composition is referred to as the active set in FW context. Many FW variants explicitly use the active set to make progress. The default variant used in Boscia, and in our set-up, is the Blended Pairwise Conditional Gradient (BPCG) method [32]. For the step-size rule, we use the secant line-search step-size rule [20], which is particularly efficient in the context of quadratic problems.
Boscia solves a Mixed-Integer Programming (MIP) problem as the LMO. By doing so, Boscia optimizes convex relaxations over the convex hull of mixed-integer feasible points instead of the continuous relaxation of the constraints, resulting in a much smaller branch-and-bound search and directly leveraging the MIP solver machinery, e.g., cutting planes, conflicts analysis, heuristics. This adjustment enables Boscia to inherently sample integer-feasible solutions, effectively embedding a heuristic search within the optimization process. Importantly, optimizing over the convex hull of integer-feasible solutions is performed by Boscia without an explicit algebraic description of this feasible region. In order to alleviate the cost of the multiple MIP solves, Boscia integrates lazified FW variants which reuse vertices of the feasible region computed by the MIP solver as much as possible, which can be performed without losing convergence guarantees; see, e.g., [12] for details on this lazification. For further lazification, Boscia holds on to vertices dropped by FW as they might become useful for nodes later down the branch. Hence, any integer feasible point is computed at most once. The active set is used during branching to facilitate warm-starting of the children nodes by splitting the active set of their parent.
MIQCQPs do not directly fit into the convex mixed-integer optimization framework that Boscia is designed for, so several key modifications to the algorithm are needed. Since Boscia assumes linear constraints, the quadratic constraints must be reformulated to comply with this structure. We achieve this by employing a power penalty relaxation, ensuring that the solver can process these constraints while maintaining feasibility, see Section 3.2.
Another key challenge is Boscia’s assumption of convexity, which does not always hold for the given MIQCQPs. To address this, we implement several modifications. First, we disable node pruning based on the lower bound in the branch-and-bound process, as the lower bound information may no longer be valid. Second, we instruct Boscia to ignore the standard FW lower bound, which is not applicable in our setting. Finally, we adjust Boscia’s solution storage mechanism. By default, it performs pre-sampling and retains only improving solutions, but this approach proved inadequate for quadratic constraints. Instead, we modify the solver to track all solutions encountered during the process, ensuring that valuable candidates are not prematurely discarded. Additionally, we added a callback mechanism that allows us to evaluate the solutions with respect to the original objective and lets us discard solutions which do not satisfy the quadratic constraints.
These modifications allow Boscia to effectively handle the challenges posed by quadratic mixed-integer problems, making it a competitive approach for the Land-Doig MIP Computational Competition 2025. Even though the focus of the competition is on the primal side, we highlight that the developed solution framework is also very suited to derive high-quality relaxation bounds, since the constraint penalization and many transformations are exact reformulations or produce relaxations of the original problem.
3 Problem transformations and presolving
In this section, we present the problem transformations performed before executing the main algorithms. These presolving steps are crucial to obtain good performance on several classes of problems.
3.1 Propagation
Initially, we apply a propagation step to tighten the bounds of the variables. For linear constraints, we apply a simple activity-based bound strengthening technique and some specialized propagators for several classes of linear constraints [16]. For quadratic constraints, we experimented with McCormick linearizations [27], introducing auxiliary variables $z_{ij}$ to replace bilinear terms $x_{i}x_{j}$, and constructing McCormick envelopes. However, we found that incorporating these linearized constraints in the problem formulation, or even just propagating them, shows little to no performance improvement in practice.
3.2 Power penalty
We take advantage of the capabilities of the Frank-Wolfe approach to solve arbitrary nonlinear objectives to relax quadratic constraints. In particular, we leverage the power penalty formulation introduced in [31] to relax quadratic constraints. Consider an objective function $f(\cdot)$ and a quadratic constraint
$x^{\top}Ax+b^{\top}x+c\leqslant 0.$
We reformulate the problem by dropping this constraint and integrating it into the objective which becomes
[…]
for a real exponent $p>1$. A parameter $p=1$ would result in a nonsmooth nondifferentiable function, while $p=2$ would result in a term akin to that of augmented Lagrangian methods. We experimented with $p\in[1.2,1.8]$, trading off steepness of the gradient towards feasible points and smoothness of the continuous subproblems tackled by Frank-Wolfe.
3.3 Quadratic special structures
In addition to the power penalty reformulation, we also exploit the specific structure of two types of quadratic constraints: complementarities and perspective constraints.
Complementarities
are constraints of the form $x_{i}x_{j}=0$. Although they are quadratic equalities, they are combinatorial in nature and can be treated as such with an additional binary variable $z$ controlling which of the two terms should be set to zero. In order to avoid relying on variable bounds for a big-M constraint, we reformulate the complementarity to a pair of indicator constraints, which we present here with the assumption that both variables are nonnegative:
$z=0\Rightarrow x_{i}\leqslant 0$
$z=1\Rightarrow x_{j}\leqslant 0.$
Importantly, we do not reformulate complementarity quadratic constraints to Special Ordered Sets of type 1 (SOS1) constraints, since such constraint would not respect the assumptions from Boscia (see [22]).
Perspective constraints
are another type of special quadratic constraint from three variables: $x$ is a single variable, $w$ is the epigraph variable, and $z$ is a binary variable. The constraint is of the form:
$x^{2}\leqslant zw\quad\text{where}\quad x\geqslant 0,w\geqslant 0,z\in\{0,1\}.$
A perspective constraint forms the convex hull of the set:
$\left\{(x,w,z)\in\mathrm{R}_{+}\times\mathrm{R}_{+}\times\{0,1\}\,:\,x^{2}\leqslant w,(1-z)x=0\right\},$
see, e.g., the seminal paper [17] on perspective functions in mixed-integer convex optimization and [19] and references therein on handling perspective functions in a nonlinear solver. The binary variable activates the continuous variable $x$, the epigraph variable $w$ is often a consequence of a nonlinear objective term converted to a constraint. When this is the case and the epigraph $w$ indeed only appears in the perspective constraint and in the objective, we transform the constraint back into a quadratic term $cx^{2}$ in the objective along with a big-M and/or indicator constraint explicitly tying $z$ and $x$. This transformation separates the nonsmooth perspective constraint that can lead to numerical challenges in a nonlinear setting (as reported in [19] among others) into a smooth quadratic term in the objective and the combinatorial “activation” structure in the constraints.
3.4 Convexification of quadratic binary problems
A subclass of MIQCQPs are quadratic binary optimization problems, which involve objective functions of the form
$\min_{x\in\{0,1\}^{n}}\frac{1}{2}x^{\top}Qx+d^{\top}x,$ (3)
where $Q\in\mathbb{R}^{n\times n}$ is a symmetric matrix and $d\in\mathbb{R}^{n}$ is a linear term. Since all variables are binary, we can leverage the identity $x_{i}^{2}=x_{i}$ to reformulate the problem and “increase” convexity.
Formally, we replace $(Q,d)$ with $(Q_{\ell},d_{\ell})$, such that the convexified matrix $Q_{\ell}$ has a proportion $\ell$ of nonnegative eigenvalues, where $0\leqslant\ell\leqslant 1$. This is achieved through spectral decomposition of $Q$, giving its eigenvalues $\lambda_{1},\dots,\lambda_{n}$ from which we can easily obtain the index $i$ such that $Q_{\ell}=Q+\lambda_{i}I$ satisfies the desired property. Accordingly, $d$ is mapped onto $d_{\ell}=d-\lambda_{i}/2$ to preserve the original objective value at all binary points.
This partial convexification balances problem tractability and the original problem structure, making convexification a tunable process controlled by the parameter $\ell$. A whole stream of work studies hardness [30] and algorithms [26] for quadratic optimization when a small number of eigenvalues are negative. Given the finite time limit, there is indeed a trade-off between the deformation of the initial objective function, whose Lipschitz constant is strongly affected by the transformation, and the advantages of mitigating nonconvexity. In practice, we explore different parameters $\ell$ in parallel on multiple threads. Computational experiments (see Section 5.6) show that some convexification indeed helps but also confirm that full convexification can be detrimental to the search for primal solutions since the original objective is dominated by the convexifying term, sometimes by orders of magnitude, .
Note that a much tighter convexification can be obtained via solving an SDP. To this end, we define a family of objective functions which are equivalent for all $x\in\{0,1\}^{n}$ via a parameter $u\in\mathbb{R}^{n}$:
$\frac{1}{2}x^{\top}(Q-2\operatorname{diag}(u))x+(d+u)^{\top}x.$
The tightest possible convexification can be obtained then via the SDP that finds the optimal vector $u$ such that $Q-2\operatorname{diag}(u)$ is positive semidefinite while minimizing the linear term. This can be formulated as:
$\min_{r,u}\;$ $r$
s.t. $\begin{bmatrix}r&-(d+u)^{\top}/2\\ -(d+u)/2&\operatorname{diag}(u)-Q/2\end{bmatrix}\succeq 0.$
We did not explore this approach though as we were primarily interested in finding primal solutions and the tightness of the convexification was not critical to the empirical performance of our approach, so that the additional computational cost of solving the SDP was not justified.
4 Heuristics
Rounding.
One of the simplest techniques is standard rounding, where fractional values are rounded to the nearest integer. Additionally, Boscia employs probability rounding for binary variables, where variables are randomly fixed to 0 or 1 based on probability distributions, and Frank-Wolfe is then used to solve for the remaining continuous variables. Another specialized rounding method is designed for 0/1 polytopes, ensuring that the rounded solution remains within the feasible region. Furthermore, for problems with simplex-like feasible regions, Boscia implements a specialized rounding heuristic that is aware of the defining hyperplane structure, improving feasibility and efficiency in these cases.
Gradient-based heuristics.
Beyond these basic rounding strategies, we utilize gradient-based heuristics, which explore a set of promising vertices, e.g., a follow-the-gradient heuristic inspired by the chasing gradients paradigm [13]. These heuristics are particularly suitable for Boscia even outside of the competition context since the algorithmic components they require (gradient and linear minimization oracles) are precisely those that are available. The method assumes an LMO-compatible feasible region; in the context of the competition, this means in particular that nonlinear constraints have been transformed to penalties. The heuristic consists in starting from any point in the feasible region, computed for instance from the LMO with an arbitrary direction. From that point, the heuristic will compute the gradient, and call the LMO to compute another extreme point. The algorithm iterates for a certain budget of maximum iterations or until it cycles to a vertex already encountered. Follow-the-gradient can be viewed as a (nonconvergent) Frank-Wolfe algorithm with a unit step size. Interestingly, the approximation guarantees of our follow-the-gradient heuristics matches the lower bounds of [1]; see also [23] for a proximity result that provides good intuition why such a heuristic can be often powerful. Preliminary computational experiments highlighted that gradient-based heuristics are not essential to performance of our overall framework; we therefore not report results for this heuristic in the main text. The reason is likely that the solutions it obtains are redundant with the ones found during the execution of Frank-Wolfe. This heuristic could however prove useful when using Boscia as an exact solver instead, with the advantage of only assuming access to the constraints as a linear minimization oracle.
Large Neighborhood Search.
Moreover, we combine our approach with the following large neighborhood search heuristics.
  1. 1.
    We introduce ASENS (Active Set Enforced Neighborhood Search), an adaptation of the RENS heuristic [3] that leverages integer vertices sampled during the execution of Frank-Wolfe algorithms. The active set refers to the set of extreme points of the feasible region whose convex combination forms the current iterate. Note that in recent variants of Frank-Wolfe, the active set size is kept small by performing correction steps and favoring steps that are local to the active set (see [33] for the blended pairwise conditional gradients variant and [7] for explanations and convergence guarantees when favoring local steps). In our case, ASENS is used only when more than 50% of variables have the same values across all points in the active set. ASENS then fixes these variables to these values and solves the subproblem. In most cases where ASENS is employed, the subproblem fixes significantly more than 50% of all variables, and the domains of the continuous variables are also reduced to the convex hull of the active set.
  2. 2.
    We implement the undercover heuristic [5], specifically designed for nonlinear problems. Undercover (UC) identifies and fixes a subset of variables such that the resulting subproblem becomes linear and in many cases easier to solve. Our implementation solves a minimum vertex cover integer linear program to determine this subset, then fixes these variables to values from a reference solution obtained from Boscia’s branch-and-bound tree. We also experimented with undercover started from multiple covers that were different enough from each other but did not observe performance improvements.
  3. 3.
    We also incorporate the solution improvement heuristic RINS (Relaxation Induced Neighborhood Search) [14] improvement heuristic, which identifies promising search regions by fixing all variables that have identical values in both the incumbent solution and the current fractional solution of the branch-and-bound tree. Also, RINS is used only when more than 50% of variables have the same values across the incumbent solution and the current fractional solution.
Specialized heuristics for QUBOs.
For QUBOs whose objective matrix has a bipartite underlying graph, an initial solution can be improved on by alternately optimizing over each component of the graph. This heuristic can seem redundant in comparison with the more general approach described above. However, the implementation of this specific case can be made very fast by exploiting this bipartite structure at a very low level in the code. The inspiration for this fast implementation comes from similar problems encountered in quantum communications, namely, the computation of the local bound of a (bipartite) Bell inequality [15]. In practice, however, this approach was overwhelmingly outperformed by the other heuristics described above.
5 Computational experiments
In this section, we want to evaluate the performance of our heuristic and the impact of its different components. We want to answer the following questions:
  • How does our heuristic perform on a diverse set of MIQCQPs?
  • What is the impact of parallelization on solution quality and efficiency?
  • How do the different LNS heuristics contribute to the overall performance?
  • What are the effects of varying the number of Frank-Wolfe iterations, power penalty parameter $p$, and convexification parameter $\ell$?
We measure performance with different metrics, including the gap at the end of the time limit (compared to a reference dual bound), primal integral (PI), and whether a solution was found altogether. These metrics highlight different important criteria that can be relevant depending on the context.
5.1 Setup
Our approach is implemented in Julia 1.11.4 on top of the Boscia and Frank-Wolfe packages. As underlying MIP solver, we use Gurobi 12.0.1. For our computational experiments, we used an Intel(R) Xeon(R) Gold 5122 @3.6GHz. We used 8 threads and a memory limit of 96GB RAM. The time limit was set to 300 seconds. We test our method on the 319 instances of the QPLIB benchmark set [18] with discrete variables. The instances are available at https://qplib.zib.de.
Evaluation Metrics
To measure the performance of our heuristic, we employ standard metrics, such as the time to first feasible solution, primal gap, and primal integral [4]. The time to first feasible solution (TTF) is the time taken to find the first feasible solution during the search process. The primal gap (Gap) of a heuristic solution $\tilde{x}$ with respect to the best known solution $x^{*}$ is defined as
$\gamma(\tilde{x})=\begin{cases}0&\text{if }|\tilde{x}|=|x^{*}|=0,\\ 1&\text{if }\tilde{x}\cdot x^{*}<0\text{ or no solution is found},\\ \dfrac{|\tilde{x}-x^{*}|}{\max(|\tilde{x}|,|x^{*}|)}&\text{otherwise}.\end{cases}$
The primal integral (PI) captures the evolution of the primal gap over time. Let $t_{0}=0$ be the start of the search process, $t_{1},\ldots,t_{s-1}$ be the points in time when a new incumbent solution is found, and $t_{s}=T$ be the end of the search process, then the primal integral is defined as
$P(T)=\sum_{i=1}^{s}\gamma(\tilde{x}_{i})\cdot(t_{i}-t_{i-1}).$
If the heuristic does not find any feasible solution, then the primal integral is equal to the time limit $T$. To report average values for these metrics across a set of instances, we use the shifted geometric mean with a shift of 1.
Parallelization Strategy
Our parallel implementation strategy executes our algorithm using varying configurations in each thread. Specifically, we ran our heuristics with different power penalty parameters $p$ ranging from 1.2 to 1.8. For problems with binary quadratic objectives, we employed various convexification strategies to explore diverse solution approaches simultaneously. The choice of different values for $p$ and convexification parameters is motivated by preliminary experiments that indicate there is no single best value for these parameters across all instances (see Section 5.6) and trying different values allows us to explore different regions of the solution space.
To further diversify the search and avoid getting trapped in local optima, we implemented a restart strategy that reinitializes our algorithm every 10 to 1000 nodes, depending on the problem characteristics. Each restart either uses the current best solution as a warm start or initiates the search with a random gradient direction to explore different regions of the solution space.
5.2 Results
In Table 1 we report our finding on QPLIB and the two categories of instances: MIQCQP and MIQP (without quadratic constraints). For each category we report the following metrics: the number of instances for which we found a feasible solution, and for instances where we found a solution, we report the average time to the first feasible solution, the optimality gap compared to the best known solution value from QPLIB, and the primal integral (PI) also using the best known solution as a reference.
Table 1: Performance of the parallel heuristic on QPLIB instances.
Category Found TTF Gap PI
MIQP 163/166 1.76 4.06 5.23
MIQCQP 98/153 7.31 25.30 50.56
All 261/319 3.18 11.57 12.77
Our heuristic is particularly effective for MIQP instances, and is able to find feasible solutions for all but three instances in this category and achieves an average optimality gap of 4.06% on instances where we found a solution. Notably, we achieved optimality (0% gap) on 104 instances. The MIQCQP category is more challenging, and we are able to find feasible solutions for 98 out of 153 instances, with an average optimality gap of 25.30%. We achieved optimality on 28 instances. In total we found feasible solutions for 261 out of 319 instances, with an average time to first feasible solution of 3.18 seconds. The average optimality gap was 11.57%, with the average primal integral at 12.77.
5.3 Impact of parallelization
Next, we compare the performance of our heuristic when run with a single master thread (base) versus when run with multiple threads (heur-parallel), with each thread exploring a different part of the search space. For a fair comparison, even when a single master thread is used, our heuristic continues to utilize all available threads for the LMO. Table 2 summarizes the results of this comparison over instances solved by at least one of the two settings, and over instances solved by both settings.
As shown in Table 2, distributing the search across multiple threads yields greater benefits than concentrating resources on the LMO calls. Specifically, the multithreaded version of our approach solves 8 more instances than the single-threaded version, reducing the average optimality gap from 16.89% to 12.07% and decreasing the average primal integral from 17.40 to 13.10, considering all instances solved by at least one of the two settings.
Over MIQPs, there is no significant difference in the number of instances solved, or the average performance metrics. However, for MIQCQPs, the multithreaded version solves 6 more instances and significantly reduces the average optimality gap and primal integral. Figure 2 shows the distribution of the optimality gap, and primal integral over MIQCQPs for both settings.
Table 2: Performance comparison between single-threaded (base) and multi-threaded (heur-parallel) approaches on QPLIB instances. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
Type Setting At Least One Solved All Solved
Found Gap (%) PI Found Gap (%) PI TTF
All (263) base 256 14.10 16.02 254 12.46 14.55 2.94
heur-parallel 261 12.07 13.10 254 11.56 12.02 2.90
MIQP (164) base 164 4.51 6.40 163 4.53 6.32 1.74
heur-parallel 163 4.47 5.38 163 4.06 5.23 1.76
MIQCQP (99) base 92 31.97 66.75 91 28.20 58.91 6.54
heur-parallel 98 25.89 51.49 91 26.37 47.73 6.24
Figure 2: Distribution of the optimality gap, and primal integral over MIQCQPs where a solution was found by at least one of the settings.
5.4 Impact of LNS heuristics
To assess the impact of the LNS heuristics finding feasible solutions, we compare the performance of the base heuristic with the LNS-off variant, which does not use any LNS heuristics, the ASENS-only variant, which only uses the ASENS heuristic, and the UC-only variant, which only uses the undercover heuristic. The results are summarized in Table 3 and demonstrate both the significant contribution of the LNS heuristics to our framework’s performance and the robustness of the underlying approach without them.
In particular, the LNS-off setting, which relies exclusively on LMO calls without any quadratic subproblem solving, successfully finds feasible solutions for a substantial amount of 226 instances, with an average optimality gap of 28.67% and a primal integral of 29.42 over the 258 instances where at least one of the settings found a solution. This result demonstrates that our framework’s core mechanism of leveraging Frank-Wolfe iterations over the convex hull of integer-feasible points is robust even when restricted to LMO calls. Similarly, the UC-only setting performs well, finding feasible solutions for 249 instances, also without solving any quadratic subproblems. The ASENS-only setting leads to an increased number of feasible solutions found, when compared to LNS-off, and achieves the closest performance with respect to the primal integral and optimality gap to the base heuristic. As shown in Table 3 and more detailed in Figure 3, the contribution of the LNS heuristics is particularly evident in the MIQCQP category, where the base heuristic finds feasible solutions for 92 instances, while the LNS-off setting finds solutions for only 74 instances, and the ASENS-only and UC-only settings find solutions for 86 and 85 instances, respectively.
Table 3: Impact of different LNS heuristics on the performance across QPLIB instances. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
Type Setting At Least One Solved All Solved
Found Gap (%) PI Found Gap (%) PI TTF
All (258) base 256 12.86 15.10 225 12.29 12.29 2.24
LNS-off 228 28.67 29.42 225 21.22 21.26 2.13
ASENS-only 243 18.66 16.53 225 14.12 12.21 2.02
UC-only 249 24.34 26.44 225 20.44 20.50 2.05
MIQP (164) base 164 4.51 6.40 153 3.65 5.29 1.44
LNS-off 154 14.56 11.82 153 10.56 9.42 1.45
ASENS-only 157 7.79 6.64 153 4.56 5.14 1.43
UC-only 164 10.97 10.74 153 10.46 9.23 1.39
MIQCQP (94) base 92 29.08 61.58 72 33.10 64.23 4.93
LNS-off 74 57.92 138.65 72 47.40 110.62 4.27
ASENS-only 86 40.16 72.93 72 37.46 66.32 3.83
UC-only 85 51.65 119.61 72 44.72 103.18 4.12
Finally, our base heuristic also includes RINS. To measure its impact, we compare the base heuristic (with RINS) against a RINS-off variant. Note that RINS, as an improvement heuristic, focuses on refining existing solutions. The impact of RINS on MIQPs is moderate since the base heuristic already finds good quality solutions for most instances. In particular, removing RINS increases the average gap and primal integral by only around 0.5%. However, for MIQCQPs, RINS provides improvements in solution quality as shown in Figure 4. Specifically, including RINS reduces the average optimality gap by 5.8% and decreases the average primal integral by a value of 11.5.
Figure 3: Distribution of the optimality gap, and primal integral over MIQCQPs where a solution was found by at least one of the settings.
Figure 4: Impact of the RINS heuristic on the performance of the base heuristic across MIQCQPs where a solution was found by at least one of the settings.
5.5 Impact of the Frank-Wolfe accuracy
We now assess the impact of the Frank-Wolfe accuracy on the performance of our heuristic, specifically by varying the number of maximum Frank-Wolfe iterations. The results for this experiment are summarized in Table 4.
Table 4: Performance of the base heuristic with different limits on the number of Frank-Wolfe iterations. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
Type Setting At Least One Solved All Solved
Found Gap (%) PI Found Gap (%) PI TTF
All (262) fw-iter=1 257 17.80 24.63 249 16.09 22.23 2.79
fw-iter=10 256 13.85 15.84 249 12.64 14.12 2.74
fw-iter=100 253 16.65 17.20 249 14.49 15.26 2.89
fw-iter=1000 253 17.50 19.78 249 15.40 17.52 3.13
MIQP (164) fw-iter=1 164 8.22 12.79 164 8.22 12.79 1.81
fw-iter=10 164 4.51 6.40 164 4.51 6.40 1.79
fw-iter=100 164 6.90 7.21 164 6.90 7.21 1.80
fw-iter=1000 164 7.99 8.12 164 7.99 8.12 1.85
MIQCQP (98) fw-iter=1 93 35.77 71.33 85 32.93 62.54 5.75
fw-iter=10 92 31.41 65.72 85 30.17 59.04 5.58
fw-iter=100 89 35.00 68.00 85 30.70 59.82 6.32
fw-iter=1000 89 35.34 81.45 85 31.16 71.63 7.46
We aim to strike a balance between computational effort in the convex relaxations and solution accuracy. Too few iterations would cause the relaxation solutions not to move sufficiently from one node to the next, while too many iterations would make the framework spend too much time computing high-accuracy solutions, which will in any case only serve as starting point to compute integer-feasible solutions through e.g., rounding afterwards. Our experiments show that 10 to 100 FW iterations per node yields a good balance between accuracy and relaxation cost, in terms of number of instances with at least one solution, mean residual gap, and primal integral.
5.6 Impact of convexification and power penalty reformulation parameters.
For the 124 MIQPs with binary variables, we tested various convexification parameters. Table 5 shows that there is no single best convexification parameter for all instances. Interestingly, fully convexifying the objective function leads to a significant increase in the average gap and primal integral.
Table 5: Performance of the base heuristic with different convexification parameters $\ell$. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
Type Setting All Solved
Found Gap (%) PI TTF
Binary QP (124) $\ell=0.6$ 124 3.66 5.17 1.46
$\ell=0.7$ 124 3.35 5.09 1.45
$\ell=0.8$ 124 3.18 4.35 1.42
$\ell=0.9$ 124 3.18 4.79 1.46
$\ell=1.0$ 124 5.02 6.62 1.41
Further we evaluated the impact of the power penalty parameter $p$ on the performance of the heuristic. The results are summarized in Table 6. We observe that the performance of the heuristic is not significantly affected by the choice of $p$ either.
Table 6: Performance of the base heuristic with different power penalty parameters $p$. Bold values are the best ones: highest number of instances solved, lowest averaged optimality gap, primal integral, and time to first feasible solution.
Type Setting At Least One Solved All Solved
Found Gap (%) PI Found Gap (%) PI TTF
MIQCQP (96) p=1.2 88 33.59 65.19 82 30.67 53.42 4.93
p=1.3 90 32.58 66.52 82 30.91 55.78 5.23
p=1.4 86 35.32 68.04 82 29.82 55.29 5.12
p=1.5 89 33.90 65.92 82 30.81 56.38 5.52
p=1.6 91 31.34 64.63 82 30.82 57.50 5.14
p=1.7 88 33.72 68.70 82 31.32 57.07 5.11
p=1.8 86 37.01 69.11 82 31.18 55.60 5.00
These observations motivate our choice of parallel and restart strategies: diversifying the search to different values of $p$ and $\ell$ maximizes the chance of finding values suited to each instance. We note that there should exist criteria to enlighten this choice based on the instance characteristics, but we leave this for further research.
5.7 Competition and Improvements on QPLIB
Our heuristic achieved first place in the Land-Doig MIP Computational Competition 2025 https://www.mixedinteger.org/2025/competition, demonstrating its effectiveness on challenging MIQCQPs. Furthermore, the heuristic found eight new best-known solutions for QPLIB instances, which are listed in Table 7 and can be found at https://qplib.zib.de/. We highlight that this improvement was achieved within the five-minute time limit of the competition.
Table 7: Improved solutions for QPLIB instances.
Instance Obj. Sense   Previous Best   New Solution    Gap
QPLIB_2169 max 29.0 30.0 3.45
QPLIB_2174 max 150.0 152.0 1.33
QPLIB_2205 max 88.0 90.0 2.27
QPLIB_3347 min 3,819,920 3,818,879 0.03
QPLIB_3584 min -25,254 -25,386 0.52
QPLIB_3709 min 5,726,530 5,710,645 0.28
QPLIB_3860 min -19,685 -20,161 2.42
QPLIB_10022 min 6,267,782 1,374,066 78.07
6 Conclusion
In this paper, we presented a primal heuristic framework for solving mixed-integer quadratically constrained quadratic programs. Our approach builds upon the Frank-Wolfe-based branch-and-bound framework Boscia and is designed to efficiently explore the solution space through gradient-guided directions and large neighborhood search heuristics that exploit integer-feasible vertices sampled during Frank-Wolfe iterations. The framework achieved first place in the Land-Doig MIP Computational Competition 2025 and discovered eight new best-known solutions for QPLIB instances within the five-minute time limit of the competition, demonstrating its practical effectiveness on challenging nonconvex mixed-integer optimization problems.